OpenStudy (18jonea):

For the geometric sequence 2, –10, 50, –250, ..., find the seventh term. –156, 250 156, 250 –31, 250 31, 250 Im not sure what to do @tkhunny

OpenStudy (18jonea):

My guess would be d but im not sure

OpenStudy (18jonea):

@tkhunny

OpenStudy (18jonea):

is x5 then -5

OpenStudy (18jonea):

@tkhunny

OpenStudy (tkhunny):

How can you be "not sure"? Count them. Term1: 2 Term 2: –10 = 2 * (-5) Term 3: 50 = 2 * (-5)^2 Term 4: –250 = 2 * (-5)^3 ... Term 7: –250 = 2 * (-5)^6 That's all it is.

OpenStudy (18jonea):

then d would be right

OpenStudy (tkhunny):

I never answer that question. YOU tell ME the correct result.

OpenStudy (18jonea):

@xapproachesinfinity

OpenStudy (18jonea):

@sweetburger

OpenStudy (18jonea):

@Mehek14

OpenStudy (18jonea):

it is d for sure

OpenStudy (tkhunny):

Seriously? Just show your intermediate results and PROVE it. You do NOT need anyone else's approval. Part of the learning of mathematics is establishing your personal confidence. Start now.

OpenStudy (tkhunny):

That's the spirit!